### v3

Abstract
Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks. However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality. In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks which is specifically designed for enhancing instruction tuning of Code Large Language Models (LLMs) and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code. Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks, which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model. Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning. Our experiments demonstrate that WaveCoder outperforms other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale. Moreover, WaveCoder exhibits high efficiency in previous code generation tasks. This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
1 Introduction
In recent months, numerous Large Language Models (LLMs) such as ChatGPT, GPT-4 [1], and Gemini 111https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks. These models utilize a self-supervised pre-training process, and subsequent fine-tuning, to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across an array of diverse tasks.
However, the overhead of holding such a large model is huge. Therefore, some relatively small LLM, especially Code Large Language Models (Code LLMs), have attracted considerable attention due to their exceptional performance on a wide range of code-related tasks. Given that LLMs can accrue a wealth of professional knowledge through pre-training [1, 2, 3], an efficient and effective pre-training process on a code corpus is critical for Code LLMs. Several previous works, including Codex [4], CodeGen [5], StarCoder [6], and CodeLLaMa [7], have successfully demonstrated that this process can significantly improve their capability to tackle code-related problems.
Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10]. These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks. InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning. Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process. Further improvements to the effectiveness of instruction tuning have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied. These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
Based on these work, our intuition is that instruction-tuning can activate the potential of LLMs then align the pre-trained models to an outstanding intelligent level. Therefore, we have summarized the main function of instruction tuning:
Generalization. Initially proposed to augment the cross-task generalization capabilities of Language Models (LMs) [8, 9, 10], instruction-tuning enhances performance across a multitude of unseen tasks when fine-tuned with diverse NLP task instructions.
Alignment. Pre-trained models [2, 3, 15, 16], having learned from a vast number of self-supervised tasks at both token and sentence levels, already possess the capability to comprehend text inputs. Instruction-tuning offers instruction-level tasks for these pre-trained models[15, 17], enabling them to extract more information from the instruction beyond the semantics of raw text. This additional information is user intention and enhances their interactive abilities with human users, thereby contributing to alignment.
To enhance the performance of Code LLMs through instruction tuning, many existing methods for generating instructional data have been devised, focusing primarily on the two aforementioned aspects. For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation. However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20]. Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs. Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
To solve these problems, as shown in Figure 2, we propose a method that could make full use of source code and explicitly control the quality of generated data. Owing to the fact that instruction tuning is to align the pre-training model to the instruction-follow training set [1, 17], we prensent a LLM Generator-Disciminator framework for instruction data generation. By employing generation and discrimination, our method can make the data generation process more customizable and more controllable. Taking the raw code as input and selecting the core dataset, our method could stably generate more realistic instruction data and control the diversity of data by adjusting the distribution of the raw code.
In this paper, our focus is on enhancing the performance of code LLMs by applying instruction-tuning. Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks: Code Summarization, Code Generation, Code Translation, and Code Repair and generate a dataset of 20,000 instruction instances, termed CodeOcean, for the four code-related tasks with our data generation strategy.
To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models, we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate them on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
2 CodeOcean: Four-task Code-related Instruction Data
In this section, we introduce the methodological details of our exploration. We firstly choose 4 representative coding tasks and collect raw code from open source dataset. For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning. The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
2.1 Tasks Setting
Given the breadth of code-related tasks from CodeXGLUE [25], we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration. These include Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
Task | Percentage(%) |
Python | 29.44 |
PHP | 21.34 |
Go | 19.68 |
Java | 18.53 |
JavaScript | 5.56 |
Others (Ruby,C++,C#) | 5.45 |
Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the model’s response is reformulated into an instructional format.
Code Generation (text-to-code, code-to-code). In this task, the model is expected to generate code based on a user’s demand description. The model generates instructions to simulate the user’s demand discourse based on the raw code. The generated raw code is then considered as the output.
Code Translation (code-to-code). This task involves converting one programming language into another. A prompt and raw code are given to the model, and it generates instructions and the translated code.
Code Repair (code-to-code). The aim of this task is to provide correct code based on potential issues in the given code. The model is expected to generate solutions for the raw code, typically within one or two sentences, which are then taken as the output.
2.2 Generation of Training Data
In past research work [26, 27], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned model. Therefore, to ensure the data quality and diversity of instruction instance, we propose: 1) a method that can retain the diversity of raw code to the utmost extent. 2) a LLM-based Generatior-Discriminator framework to further improve the data quality of instruction instances.
2.2.1 Raw Code Collection
To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy [28, 29] to get the raw code collection from the open source code dataset for instruction data generation. In this work, we select CodeSearchNet 222https://huggingface.co/datasets/code_search_net which contains 2 million (comment, code) pairs from open-source libraries hosted on GitHub and includes code and documentation for 6 programming languages as our foundation dataset and develop a series of manually filtering rules to ensure the basic data quality of raw code. Subsequently, we apply the coreset-based selection method KCenterGreedy to maximize the diversity of the raw code. More details are as follows:
Manually defined filtering rules. In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
i)
The length of required code should not be too long or too short. In this work, we filtered the code whose length lower than 50 or higher than 800.
ii)
Followed Code Alpaca333https://github.com/sahil280114/codealpaca, we have eliminated the raw code containing words from the blacklist, since these code could potentially reduce the performance of the resulting model.
Coreset-based selection method. After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune. To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
In this method, our main work is divided into two steps:
i)
Code embedding. In this step, we encode all the raw code samples using a pre-trained language model (roberta-large-v1 [16] in this paper). we take the [CLS] token as the code embedding for one input raw code, which is not defined in generative language models like GPT-3.5.
ii)
Coreset sampling. In the coreset sampling step, our goal is to use a small number of samples to represent the distribution of the overall data, thus we employed unsupervised clustering and focused on the data representations. After obtaining the code embeddings from the first step, we used KCentergreedy algorithm [28] to select the core dataset and maximize data diversity, which aims to choose center points such that minimize the largest distance between a random data point and its nearest center, and has been proven efficient in obtaining a set of core samples of one distribution.
Task | Num | Per(%) | Prompt |
Code Generation | 11370 | 57.0 | Implementing functions that perform specific operations given input. |
Code Summarization | 3175 | 15.9 | Write clear and concise documentation for the given code. |
Code Repair | 3144 | 15.7 | Identify and fix errors in the given code. |
Code Translation | 2236 | 11.2 | Rewrite the given code from one programming language to another. |
2.2.2 LLM-based Generator-Discriminator Framework
After the process of collecting raw code, it is more significant to generate high-quality and diverse instruction data and reformat them to training data. To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework, where the Generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data. and the Discriminator can continue analyzing and filtering the instruction data. By incorporating such a framework into the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself and the data quality of instruction data generation can be enhanced by the Discriminator.
Generation Phase. In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario. As shown in Figure 3, following the model-generated task definition, we manually develop the generation requirements for the each code-related task. Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select good and bad case examples from the example database to leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning. However, based on our experiments, we observe that few-shot generation is not always the best way to get instruction data especially on code generation task. We analyzed that this is because the teacher LLM always imitated the given example during the few-shot generation process, which will damage the diversity of instruction data. Therefore, we employed zero-shot generation on code generation task and few-shot generation on other tasks.
Discrimination Phase. During the exploration of the instruction generation process, we noticed that the data quality of the instruction instances cannot be guaranteed through the generation phase alone. In order to enhance the controllability of data generation and further ensure the data quality, we designed a LLM-based discriminator to continue analyzing and filtering the instruction data. For the discrimination phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances. However, due to the large number of rules, the discriminator cannot conduct effective and comprehensive discrimination. To solve this problem, we disassemble the discrimination rules to some subtopics inspired by Zero-shot-CoT [30] and employ the LLM-based discriminator to analyze the generation step by step . By adopting this method, the discrimination process can be more contrallable when filtering low-quality instruction instances. After the discrimination process, each instruction instance is classified as either a good or bad case and the classification information is subsequently reused in the following generation as examples. For the reusage of these classified instruction instance, different from self-instruct [31] which solely utilize the exemplification as good example, we exploit both the good generation and bad generation as the few-shot example so that the generator can gain from the analysis of bad example. An example of our Generator and Discriminator framework is presented in Figure 4. This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
3 Experiments
3.1 Training
Unlike the previous work [14, 32, 33] that mainly focus on code generation task, we hope that CodeOcean could gain more general problem-solving capabilities on a wider range of coding tasks. Therefore, our training relies on a 20k instruction dataset generated by our LLM-based Generator-Discriminator framework and an additional logical dataset. The 20k dataset covers 4 common code tasks, including code generation, code translation, code summarization, and code repair.
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. Followed Alpaca [12], we set the fine-tuning prompt as follows:
3.2 Evaluation
Model | Params | Base Model | InsT Data | HumanEval | MBPP (500) |
Proprietary Models | |||||
GPT-4 [1] | - | - | - | 86.6 / 67.0 | - |
ChatGPT [1] | - | - | - | 73.2 / 48.1 | 52.2 |
Open-Source Models | |||||
StarCoder | 15B | - | ✘ | 33.6 | 43.3 |
OctoCoder [21] | 15B | StarCoder | 13K | 46.2 | 43.5 |
WizardCoder [14] | 15B | StarCoder | 78K | 57.3 | 51.8 |
WaveCoder-SC-15B ![]() |
15B | StarCoder | 20K | 50.5 (+16.9) | 51.0 (+7.4) |
CodeLLaMa [7] | 7B | - | ✘ | 33.5 | 41.4 |
CodeLLaMa-instruct [7] | 7B | CodeLLaMa | 14K | 34.8 | 44.4 |
WaveCoder-CL-7B ![]() |
7B | CodeLLaMa | 20K | 48.1 (+14.6) | 47.2 (+5.8) |
CodeLLaMa [7] | 13B | - | ✘ | 36.0 | 47.0 |
CodeLLaMa-instruct [7] | 13B | CodeLLaMa | 14K | 42.5 | 49.4 |
WaveCoder-CL-13B ![]() |
13B | CodeLLaMa | 20K | 55.4 (+19.4) | 49.6 (+2.6) |
DeepseekCoder [22] | 6.7B | - | ✘ | 49.4 | 60.6 |
Magicoder-DS[34] | 6.7B | DeepseekCoder | 75K | 66.5 | 60.4 |
WaveCoder-DS-6.7B ![]() |
6.7B | DeepseekCoder | 20K | 64.0 (+14.6) | 62.8 (+2.2) |
3.2.1 Benchmarks and Baselines
HumanEval 444https://huggingface.co/datasets/openai_humaneval, which consists of 164 manually-written Python programming problems and an average of 9.6 test cases allocated to each problem is now the most extensively adopted benchmark for Code LLMs.
MBPP 555https://huggingface.co/datasets/mbpp consists of around 1,000 crowd-sourced Python programming problems, designed to be solvable by entry level programmers, covering programming fundamentals, standard library functionality, and so on. In this paper, we choose the 500 problems test dataset to evaluate both few-shot inference of fine-tuned models. For whose MBPP (500) result is not reported or not used, we reproduced for them using bigcode-evaluation-harness 666https://github.com/bigcode-project/bigcode-evaluation-harness.
HumanEvalPack 777https://huggingface.co/datasets/bigcode/humanevalpack is an extension of OpenAI’s HumanEval to cover 6 total languages across 3 tasks. In this paper, we select the HumanEvalFix to evaluate the code to code ability especially on code repair task and HumanEvalExplain benchmarks to evaluate the code to text ability especially on code summarization task.
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack [21] or evaluate by ourselves.
Open Source Models. Base models comprise StarCoder, CodeLLaMa and Deepseek-Coder with prompting. To ensure an equitable comparison, we opted to select those instructed models that have been trained using fewer than 100,000 instruction instances for our comparative analysis.
3.2.2 metrics
Pass@k. Followed most previous works, we also adopt the pass@k metric proposed by OpenAI [4] to evaluate the performance of our model. By generating samples for each promblem and checking the correctness for each sample with the test cases, the pass@k score can be calculated by the unbiased estimator:
where is the total number of sample generated for a problem, is the number of correct samples passing all the test cases.
3.2.3 Evaluation on Code Generation Task
HumanEval and MBPP are two representative benchmarks for code generation task where the model should generate the complete code based on the function signature and the docstring of the problem. Table 3 shows the pass@1 score of different LLMs on both benchmarks. From the results, We have the following observations:
WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data) and CodeAlpaca with 20K data. Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. According to the comparison in Table 3, WaveCoder solely relies on a dataset encompassing 20,000 entries to accomplish performance levels on par with those of Magicoder [34] and Wizardcoder [14]. Although WaveCoder falls behind WizardCoder [14] and Magicoder [34] on code generation task, there is a considerable difference in data size among them. Nevertheless, the performance lag of WaveCoder is relatively minor.
Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
GPT-4 | 47.0 | 48.2 | 50.0 | 50.6 | 47.6 | 43.3 | 47.8 |
StarCoder | 8.7 | 15.7 | 13.3 | 20.1 | 15.6 | 6.7 | 13.4 |
OctoCoder | 30.4 | 28.4 | 30.6 | 30.2 | 26.1 | 16.5 | 27.0 |
WizardCoder | 31.8 | 29.5 | 30.7 | 30.4 | 18.7 | 13.0 | 25.7 |
WaveCoder-SC-15B ![]() |
39.3 | 35.1 | 34.8 | 36.2 | 30.2 | 22.5 | 33.0 |
CodeLLaMa-instruct-7B | 28.0 | 23.2 | 23.2 | 18.3 | 0.1 | 0.1 | 15.5 |
CodeLLaMa-CodeAlpaca-7B | 37.8 | 39.0 | 42.0 | 37.8 | 37.2 | 29.2 | 37.1 |
WaveCoder-CL-7B ![]() |
41.4 | 41.4 | 42.0 | 47.1 | 42.7 | 34.7 | 41.5 |
CodeLLaMa-instruct-13B | 29.2 | 19.5 | 32.3 | 24.4 | 12.8 | 0.1 | 19.7 |
CodeLLaMa-CodeAlpaca-13B | 42.7 | 43.9 | 50.0 | 45.7 | 39.6 | 37.2 | 43.2 |
WaveCoder-CL-13B ![]() |
48.8 | 48.2 | 50.6 | 51.8 | 45.1 | 40.2 | 47.4 |
DeepseekCoder-6.7B | 29.9 | 29.2 | 39.0 | 29.2 | 25.0 | 21.9 | 29.0 |
Magicoder-DS | 42.0 | 43.3 | 50.6 | 41.4 | 38.4 | 29.2 | 40.8 |
DeepseekCoder-CodeAlpaca-6.7B | 49.4 | 51.8 | 45.1 | 48.8 | 44.5 | 31.7 | 45.2 |
WaveCoder-DS-6.7B ![]() |
57.9 | 52.4 | 57.3 | 47.5 | 45.1 | 36.0 | 49.4 |
3.2.4 Evaluation on Other Code-related Task
We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4, 5 and use the result reported in Octopack [21] when avaliable. Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
WaveCoder outperforms all open source models on other code-related task with similar degrees of fine-tuning. Using Starcoder as the foundation model, our WaveCoder-SC outperforms both WizardCoder and OctoCoder across all programming languages on HumanEvalFix and HumanEvalExplain benchmarks. Even when compared with other foundation models, our model consistently demonstrates superior performance over all previously fine-tuned models.
Enhancing the refinement and diversity of data can significantly improve the effectiveness of instruction tuning in multi-task scenarios. By refining the data and classifying the instructions to 4 different code-related downstream tasks, our models achieve a surprising performance level on both evaluation scenarios. Notably, WaveCoder-DS, utilizing only 6.7B parameters, approximates the performance of GPT-4 on HumanEvalFix quite closely.
Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
GPT-4 | 64.6 | 57.3 | 51.2 | 58.5 | 38.4 | 42.7 | 52.1 |
StarCoder | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
WizardCoder | 32.5 | 33.0 | 27.4 | 26.7 | 28.2 | 16.9 | 27.5 |
OctoCoder | 35.1 | 24.5 | 27.3 | 21.1 | 24.1 | 14.8 | 24.5 |
WaveCoder-SC-15B ![]() |
37.1 | 33.3 | 40.5 | 23.3 | 31.8 | 19.3 | 30.8 |
CodeLLaMa-instruct-7B | 33.5 | 36.0 | 31.7 | 21.3 | 25.0 | 16.4 | 27.3 |
CodeLLaMa-CodeAlpaca-7B | 34.7 | 24.4 | 37.8 | 23.2 | 28.6 | 19.5 | 28.0 |
WaveCoder-CL-7B ![]() |
41.4 | 31.7 | 39.0 | 25.0 | 34.1 | 23.2 | 32.4 |
CodeLLaMa-instruct-13B | 40.2 | 26.8 | 37.2 | 22.5 | 28.0 | 14.6 | 28.2 |
CodeLLaMa-CodeAlpaca-13B | 32.3 | 28.0 | 34.1 | 18.9 | 29.9 | 20.7 | 27.3 |
WaveCoder-CL-13B ![]() |
45.7 | 42.0 | 48.2 | 32.3 | 38.4 | 20.7 | 37.9 |
DeepseekCoder-6.7B | 43.9 | 40.2 | 37.8 | 29.2 | 34.1 | 22.5 | 34.6 |
Deepseek-CodeAlpaca-6.7B | 40.8 | 37.2 | 42.1 | 29.9 | 31.7 | 22.5 | 34.0 |
Magicoder-DS | 55.5 | 36.6 | 49.4 | 36.0 | 39.6 | 27.4 | 40.7 |
WaveCoder-DS-6.7B ![]() |
48.2 | 47.5 | 49.4 | 32.3 | 48.2 | 22.0 | 41.3 |
3.3 Comparison with CodeAlpaca
CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct [12]. To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeOcean), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts. To prevent overfitting, we use Low-Rank Adaption (LoRA) [35] for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset. Table 4, 5 present the result of CodeOcean on HumanEval and other code-related benchmarks highlighting the following salient observations:
1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1. Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data. As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
2) According to Table 4 and Table 5, All WaveCoder models significantly outperform the model fine-tuned with CodeAlpaca. Remarkably, The pass@1 score of WaveCoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain. This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
4 Ablation and Analysis
To evaluate the the effectiveness of each code-related task appearing in training data and the relationship between different tasks, we conduct an ablation study about the task type of instruction data. Using DeepseekCoder-Base-6.7B as our base model, we have the following observations from Table 6:
Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff. As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks. For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
Different tasks can promote each other so that the model can show a generalization ability. From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks. For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks. Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
Model | CG | CS | CT | CR | HumanEval | HumanEval Fix (Avg.) | HumanEval Explain (Avg.) |
DeepseekCoder-Base-6.7B | ✘ | ✘ | ✘ | ✘ | 49.4 | 29.0 | 34.6 |
WaveCoder-DS-6.7B ![]() |
✔ | ✔ | ✔ | ✔ | 64.0 (+14.6) | 49.4 (+20.4) | 41.3 (+7.3) |
-Without Code Repair | ✔ | ✔ | ✔ | ✘ | 60.9 (-3.1) | 15.7 (-33.7) | 41.2 (-0.1) |
-Without Code Generation | ✘ | ✔ | ✔ | ✔ | 53.6 (-10.4) | 47.4 (-2.0) | 40.5 (-0.8) |
-Without Code Translation | ✔ | ✔ | ✘ | ✔ | 60.9 (-3.1) | 49.3 (-0.1) | 41.6 (+0.3) |
-Without Code Summarization | ✔ | ✘ | ✔ | ✔ | 61.5 (-2.5) | 45.6 (-3.8) | 28.4 (-12.9) |
5 Related Work
5.1 Code Large Language Models
Numerous researchers have proposed their Code Large Language Models (LLMs), including CodeGen [36], CodeT5[37], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain. By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios. In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [38] and WizardCoder [14] have been introduced. When drafting this paper, we noticed that our contemporaneous work Magicoder [34] has been released, so we conduct a brief comparative analysis in Section 3.
5.2 Instruction Tuning
In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language. For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users. This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT. Flan-PaLM 540B [10] is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model. Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training. For Code LLMs, InstructCodeT5+ [38], WizardCoder [14], Pangu-coder2 [32], are recent recent models trained with coding instructions. However, all the instruction data they used is from CodeAlpaca, which is not refined enough.
6 Conclusion
This paper studies the effect of multi-task instruction data on enhancing the generalization ability of Code Large Language Model. We introduce CodeOcean, a dataset with 20K instruction instances on four universal code-related tasks. This method and dataset enable WaveCoder, which significantly imporove the generalization ability of foundation model on diverse downstream tasks. Our WaveCoder has shown the best generalization ability among other open-source models in code repair and code summarization tasks, and can maintain high efficiency on previous code generation benchmarks. In order to further improve the mono-task performance and generalization ability of the model, future work should focus on the interplay among different tasks and larger dataset.

### v4

@@
-
+WaveCoder: Widespread And Versatile Enhancement For Code Large Language Models By Instruction Tuning
 Abstract
-Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks.
-However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality.
-In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks which is specifically designed for enhancing instruction tuning of Code Large Language Models (LLMs) and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code.
-Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks, which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model.
-Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning.
-Our experiments demonstrate that WaveCoder outperforms other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale.
-Moreover, WaveCoder exhibits high efficiency in previous code generation tasks.
-This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
+Recent work demonstrates that, after instruction tuning, Code Large Language Models (Code LLMs) can obtain impressive capabilities to address a wide range of code-related tasks.
+However, current instruction tuning methods for Code LLMs mainly focus on the traditional code generation task, resulting in poor performance in complex multi-task scenarios.
+In this paper, we concentrate on multiple code-related tasks and present WaveCoder, a series of Code LLMs trained with Widespread And Versatile Enhanced instruction data.
+To enable the models to tackle complex code-related tasks, we propose a method to stably generate diverse, high-quality instruction data from open source code dataset in multi-task scenarios and obtain CodeSeaXDataset, a dataset comprising 19,915 instruction instances across 4 code-related tasks, which is aimed at improving the generalization ability of Code LLM.
+Our experiments demonstrate that WaveCoder models significantly outperform other open-source models in terms of the generalization ability across different code-related tasks.
+Moreover, WaveCoder-Ultra-6.7B presents the state-of-the-art generalization abilities on a wide range of code-related tasks.
 1 Introduction
-In recent months, numerous Large Language Models (LLMs) such as ChatGPT, GPT-4 [1], and Gemini 111https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks.
-These models utilize a self-supervised pre-training process, and subsequent fine-tuning, to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across an array of diverse tasks.
-However, the overhead of holding such a large model is huge.
-Therefore, some relatively small LLM, especially Code Large Language Models (Code LLMs), have attracted considerable attention due to their exceptional performance on a wide range of code-related tasks.
-Given that LLMs can accrue a wealth of professional knowledge through pre-training [1, 2, 3], an efficient and effective pre-training process on a code corpus is critical for Code LLMs.
-Several previous works, including Codex [4], CodeGen [5], StarCoder [6], and CodeLLaMa [7], have successfully demonstrated that this process can significantly improve their capability to tackle code-related problems.
-Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10].
-These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks.
-InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning.
-Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process.
-Further improvements to the effectiveness of instruction tuning have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied.
-These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
-Based on these work, our intuition is that instruction-tuning can activate the potential of LLMs then align the pre-trained models to an outstanding intelligent level.
-Therefore, we have summarized the main function of instruction tuning:
-Generalization.
-Initially proposed to augment the cross-task generalization capabilities of Language Models (LMs) [8, 9, 10], instruction-tuning enhances performance across a multitude of unseen tasks when fine-tuned with diverse NLP task instructions.
-Alignment.
-Pre-trained models [2, 3, 15, 16], having learned from a vast number of self-supervised tasks at both token and sentence levels, already possess the capability to comprehend text inputs.
-Instruction-tuning offers instruction-level tasks for these pre-trained models[15, 17], enabling them to extract more information from the instruction beyond the semantics of raw text.
-This additional information is user intention and enhances their interactive abilities with human users, thereby contributing to alignment.
-To enhance the performance of Code LLMs through instruction tuning, many existing methods for generating instructional data have been devised, focusing primarily on the two aforementioned aspects.
-For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation.
-However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20].
-Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs.
-Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
-To solve these problems, as shown in Figure 2, we propose a method that could make full use of source code and explicitly control the quality of generated data.
-Owing to the fact that instruction tuning is to align the pre-training model to the instruction-follow training set [1, 17], we prensent a LLM Generator-Disciminator framework for instruction data generation.
-By employing generation and discrimination, our method can make the data generation process more customizable and more controllable.
-Taking the raw code as input and selecting the core dataset, our method could stably generate more realistic instruction data and control the diversity of data by adjusting the distribution of the raw code.
-In this paper, our focus is on enhancing the performance of code LLMs by applying instruction-tuning.
-Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks: Code Summarization, Code Generation, Code Translation, and Code Repair and generate a dataset of 20,000 instruction instances, termed CodeOcean, for the four code-related tasks with our data generation strategy.
-To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models, we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate them on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
-2 CodeOcean: Four-task Code-related Instruction Data
-In this section, we introduce the methodological details of our exploration.
-We firstly choose 4 representative coding tasks and collect raw code from open source dataset.
-For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning.
-The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
-2.1 Tasks Setting
-Given the breadth of code-related tasks from CodeXGLUE [25], we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration.
-These include Code Summarization, Code Generation, Code Translation, and Code Repair.
+Recently, Large Language Models (LLMs) such as ChatGPT, GPT-4 OpenAI (2023), and Gemini 111https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks.
+These models utilize a self-supervised pre-training process, and subsequent supervised fine-tuning to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across various tasks.
+For code-related tasks, several previous works, including Codex Chen et al.
+(2021), StarCoder Li et al.
+(2023a), CodeLLaMa Roziere et al.
+(2023) and DeepseekCoder Guo et al.
+(2024), have successfully demonstrated that pre-training on code corpus can significantly improve the model’s capability to tackle code-related problems.
+After the process of pre-training, instruction tuning Wei et al.
+(2022); Aribandi et al.
+(2022); Chung et al.
+(2022) has shown its effectiveness in the aspect of improving the quality of LLM responses.
+To specifically enhance the performance of Code LLMs on code-related tasks through instruction tuning, many existing methods for instruction data generation have been designed.
+For example, Code Alpaca Chaudhary (2023) utilizes the method of self-instruct Wang et al.
+(2023a) within the coding domain, leveraging the few-shot capabilities of teacher LLM to generate instruction data.
+Similarly, WizardCoder Luo et al.
+(2024) applies the evol-instruct Xu et al.
+(2024) approach based on Code Alpaca, demonstrating a novel and effective method for the generation of instruction data.
+These applications underscore the potential of utilizing teacher LLMs to produce instructional content effectively, thereby offering an avenue for the creation of instruction data in the code domain.
+However, the quality of the data they generate heavily relies on the performance of the teacher LLM and the limited initial seeds, which often produces a large amount of duplicate instruction instances and reduce the effectiveness of instruction tuning Xu et al.
+(2022); Yan et al.
+(2023); Lee et al.
+(2022).
+To break away from dependence on teacher LLMs, Octopack Muennighoff et al.
+(2024) constructs a code instruction dataset leveraging the natural structure of Git commits.
+Nonetheless, ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
+Additionally, these endeavors are predominantly centered on traditional code generation tasks and lack the capability to produce detailed, task-specific instructions in multi-task scenarios.
+In this paper, we primarily focus on multiple code-related tasks, aiming to generate high-quality and diverse instructional data tailored to specific task requirements.
+Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks in CodeXGLUE Lu et al.
+(2021): 1) Code Summarization, 2) Code Generation, 3) Code Translation, 4) Code Repair and propose a widespread and versatile enhanced instruction generation method that could make full use of open source code data and stably generate high quality and diverse instruction data in multi-task scenarios.
+By this generation strategy, we obtain a dataset of 19,915 instruction instances across four code-related tasks, termed CodeSeaXDataset.
+To validate our approach, we train StarCoder Li et al.
+(2023a), CodeLLaMa Roziere et al.
+(2023), and DeepseekCoder Guo et al.
+(2024) with our initial CodeSeaXDataset dataset and get WaveCoder.
+Following a thorough assessment on HumanEval Chen et al.
+(2021), MBPP Austin et al.
+(2021), HumanEvalPack Muennighoff et al.
+(2024) benchmarks, experimental results show that our WaveCoder exhibits outstanding generalization ability based on widespread and versatile enhanced instruction tuning.
+Moreover, to further explore the improvements brought by data quality, we use GPT-4 OpenAI (2023) to regenerate response for the instruction in CodeSeaXDataset.
+Fine-tuned with the enhanced 20K CodeSeaXDataset dataset, we obtain WaveCoder-Pro-6.7B which achieve 72.0% pass@1 on HumanEval Chen et al.
+(2021) and surpass open source Code LLMs but still behind SoTA Code LLM.
+Combining enhanced CodeSeaXDataset with WaveCoder-evol-codealpaca, the decontaminated Magicoder-evol-codealpaca 222https://huggingface.co/datasets/ise-uiuc/Magicoder-evol-codealpaca-110K dataset, we present WaveCoder-Ultra-6.7B, with SoTA generalization capabilities on multiple code-related tasks.
+2 CodeSeaXDataset: Four-task Code-related Instruction Data
+2.1 Tasks Details
+| Task | Num | Per(%) | Prompt |
+| Code Generation | 11370 | 57.1 | Implementing functions that perform specific operations given input. |
+| Code Summarization | 3165 | 15.8 | Write clear and concise documentation for the given code. |
+| Code Repair | 3144 | 15.8 | Identify and fix errors in the given code. |
+| Code Translation | 2236 | 11.2 | Rewrite the given code from one programming language to another. |
+Given the code-related tasks from CodeXGLUE Lu et al.
+(2021), we select four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration including Code Summarization, Code Generation, Code Translation, and Code Repair.
 Detailed descriptions of these tasks can be found below.
@@
 | Others (Ruby,C++,C#) | 5.45 |
-Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the model’s response is reformulated into an instructional format.
+Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the teacher model’s response is reformulated into an instruction format.
 Code Generation (text-to-code, code-to-code).
 In this task, the model is expected to generate code based on a user’s demand description.
-The model generates instructions to simulate the user’s demand discourse based on the raw code.
-The generated raw code is then considered as the output.
+Therefore, the teacher model is expected to generate instructions and solution code given the raw code as a instruction-solution pair.
+The generated solution code is then considered as the output.
 Code Translation (code-to-code).
 This task involves converting one programming language into another.
-A prompt and raw code are given to the model, and it generates instructions and the translated code.
+The task-specific prompt and raw code are given to the teacher model, then the model generates instructions and the translated code.
 Code Repair (code-to-code).
 The aim of this task is to provide correct code based on potential issues in the given code.
-The model is expected to generate solutions for the raw code, typically within one or two sentences, which are then taken as the output.
-2.2 Generation of Training Data
-In past research work [26, 27], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount.
-The improvement of data quality and diversity are directly related to the performance of the fine-tuned model.
-Therefore, to ensure the data quality and diversity of instruction instance, we propose: 1) a method that can retain the diversity of raw code to the utmost extent.
-2) a LLM-based Generatior-Discriminator framework to further improve the data quality of instruction instances.
+The teacher model is expected to generate solutions for the incorrect code, typically with the correct code and some descriptions, which are then taken as the output.
+2.2 Widespread and Versatile Enhanced Instruction Generation
+In past research work Zhou et al.
+(2023); Gupta et al.
+(2023), many researchers have discovered that data quality and diversity often play a more important role in instruction tuning process than data amount.
+The improvement of data quality and diversity are directly related to the performance of the fine-tuned LLM.
+Therefore, to ensure the data quality and diversity of instruction instance, we propose a widespread and versatile enhanced instruction generation method including two the following parts: 1) a method that can retain the diversity of instruction data by retainig the diversity of raw code to the utmost extent.
+2) a LLM-based Generator-Discriminator framework to stably generate high-quality instruction data.
 2.2.1 Raw Code Collection
-To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy [28, 29] to get the raw code collection from the open source code dataset for instruction data generation.
-In this work, we select CodeSearchNet 222https://huggingface.co/datasets/code_search_net which contains 2 million (comment, code) pairs from open-source libraries hosted on GitHub and includes code and documentation for 6 programming languages as our foundation dataset and develop a series of manually filtering rules to ensure the basic data quality of raw code.
-Subsequently, we apply the coreset-based selection method KCenterGreedy to maximize the diversity of the raw code.
-More details are as follows:
-Manually defined filtering rules. In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
-i)
-The length of required code should not be too long or too short. In this work, we filtered the code whose length lower than 50 or higher than 800.
-ii)
-Followed Code Alpaca333https://github.com/sahil280114/codealpaca, we have eliminated the raw code containing words from the blacklist, since these code could potentially reduce the performance of the resulting model.
-Coreset-based selection method.
-After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune.
-To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
-In this method, our main work is divided into two steps:
-i)
-Code embedding.
-In this step, we encode all the raw code samples using a pre-trained language model (roberta-large-v1 [16] in this paper).
-we take the [CLS] token as the code embedding for one input raw code, which is not defined in generative language models like GPT-3.5.
-ii)
-Coreset sampling.
-In the coreset sampling step, our goal is to use a small number of samples to represent the distribution of the overall data, thus we employed unsupervised clustering and focused on the data representations.
-After obtaining the code embeddings from the first step, we used KCentergreedy algorithm [28] to select the core dataset and maximize data diversity, which aims to choose center points such that minimize the largest distance between a random data point and its nearest center, and has been proven efficient in obtaining a set of core samples of one distribution.
-Task | Num | Per(%) | Prompt |
-Code Generation | 11370 | 57.0 | Implementing functions that perform specific operations given input. |
-Code Summarization | 3175 | 15.9 | Write clear and concise documentation for the given code. |
-Code Repair | 3144 | 15.7 | Identify and fix errors in the given code. |
-Code Translation | 2236 | 11.2 | Rewrite the given code from one programming language to another. |
+To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy Sener and Savarese (2018); Chen et al.
+(2023) to get the raw code collection from the open source code dataset.
+In this work, we select CodeSearchNet 333https://huggingface.co/datasets/code_search_net, which contains 2 million of <comment, code> pairs from open-source libraries hosted on GitHub, as our foundation dataset and process it with the following steps:
+Manually defined filtering rules.
+In order to select high-quality code for instruction-tuning, we make the following rules to filter the foundation dataset: i) In this work, we filtered the code to make sure that the length of required code is neither too long nor too short.
+ii) Followed Code Alpaca Chaudhary (2023), we have eliminated the raw code containing words from the blacklist, which could potentially reduce the performance of the resulting model.
+Coreset selection method.
+To ensure the data diversity when select raw code samples, we employed KCenterGreedy Sener and Savarese (2018) algorithm, which has been proven efficient in obtaining a set of core samples of one distribution, to select representative samples from the open source code dataset based on the code embeddings encoded by the same embedding model (roberta-large-v1 Liu et al.
+(2019)).
+By incorporating such a method into the open source code dataset, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself or initial seed.
+Moreover, due to the application of the KCenterGreedy algorithm, the diversity of languages is also significantly retained, as shown in Table 2.
+| Model | Params | Base Model | InsT Data | HumanEval | MBPP (500) |
+| Proprietary Models | |||||
+| GPT-4 | - | - | - | 85.4 / 67.0 | - |
+| ChatGPT | - | - | - | 73.2 / 48.1 | 52.2 |
+| Open-Source Models | |||||
+| StarCoder | 15B | - | ✘ | 33.6 | 43.3 |
+| OctoCoder | 15B | StarCoder | 13K | 46.2 | 43.5 |
+| WizardCoder | 15B | StarCoder | 78K | 57.3 | 51.8 |
+| WaveCoder-SC-15B |
+15B | StarCoder | 20K | 50.5 (+16.9) | 51.0 (+7.4) |
+| CodeLLaMa | 7B | - | ✘ | 33.5 | 41.4 |
+| CodeLLaMa-instruct | 7B | CodeLLaMa | 14K | 34.8 | 44.4 |
+| WaveCoder-CL-7B |
+7B | CodeLLaMa | 20K | 48.1 (+14.6) | 47.2 (+5.8) |
+| CodeLLaMa | 13B | - | ✘ | 36.0 | 47.0 |
+| CodeLLaMa-instruct | 13B | CodeLLaMa | 14K | 42.5 | 49.4 |
+| WaveCoder-CL-13B |
+13B | CodeLLaMa | 20K | 55.4 (+19.4) | 49.6 (+2.6) |
+| DeepseekCoder | 6.7B | - | ✘ | 49.4 | 60.6 |
+| Magicoder-DS | 6.7B | DeepseekCoder | 75K | 66.5 | 60.4 |
+| WaveCoder-DS-6.7B |
+6.7B | DeepseekCoder | 20K | 64.0 (+14.6) | 62.8 (+2.2) |
+| WaveCoder-Pro-6.7B |
+6.7B | DeepseekCoder | 20K | 72.0 (+22.6) | 63.6 (+3.0) |
+| SoTA Open-Source Models | |||||
+| DeepseekCoder-instruct∗ | 6.7B | DeepseekCoder | - | 73.8 | 62.8 |
+| Magicoder-S-DS | 6.7B | DeepseekCoder | 185K | 76.8 | 64.6 |
+| WaveCoder-Ultra-6.7B |
+6.7B | DeepseekCoder | 130K | 78.6 (+29.2) | 64.4 (+3.8) |
 2.2.2 LLM-based Generator-Discriminator Framework
-After the process of collecting raw code, it is more significant to generate high-quality and diverse instruction data and reformat them to training data.
-To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework, where the Generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data.
-and the Discriminator can continue analyzing and filtering the instruction data.
-By incorporating such a framework into the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself and the data quality of instruction data generation can be enhanced by the Discriminator.
+After the process of raw code collection, the data diversity from raw code has been retained, where the next step is to generate instruction data for supervised fine-tuning from the raw code.
+To further ensure the quality of generated instruction data, shown in Figure 2, we propose a LLM-based Generator-Discriminator framework where the generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data and the discriminator can generate analysis for each component in instruction data.
 Generation Phase.
-In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario.
+In the generation phase, we utilize GPT-4 to generate definitions for each code-related task.
 As shown in Figure 2, following the model-generated task definition, we manually develop the generation requirements for the each code-related task.
-Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select good and bad case examples from the example database to leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning.
-However, based on our experiments, we observe that few-shot generation is not always the best way to get instruction data especially on code generation task.
-We analyzed that this is because the teacher LLM always imitated the given example during the few-shot generation process, which will damage the diversity of instruction data.
-Therefore, we employed zero-shot generation on code generation task and few-shot generation on other tasks.
+Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select different examples from the example database to generate instruction data by GPT-3.5.
 Discrimination Phase.
 During the exploration of the instruction generation process, we noticed that the data quality of the instruction instances cannot be guaranteed through the generation phase alone.
-In order to enhance the controllability of data generation and further ensure the data quality, we designed a LLM-based discriminator to continue analyzing and filtering the instruction data.
-For the discrimination phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances.
-However, due to the large number of rules, the discriminator cannot conduct effective and comprehensive discrimination.
-To solve this problem, we disassemble the discrimination rules to some subtopics inspired by Zero-shot-CoT [30] and employ the LLM-based discriminator to analyze the generation step by step .
-By adopting this method, the discrimination process can be more contrallable when filtering low-quality instruction instances.
-After the discrimination process, each instruction instance is classified as either a good or bad case and the classification information is subsequently reused in the following generation as examples.
-For the reusage of these classified instruction instance, different from self-instruct [31] which solely utilize the exemplification as good example, we exploit both the good generation and bad generation as the few-shot example so that the generator can gain from the analysis of bad example.
-An example of our Generator and Discriminator framework is presented in Figure 4.
-This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
+In order to enhance the controllability of data generation and further ensure the data quality, we employ GPT-4 as a LLM-based discriminator to continue analyzing and filtering the instruction data.
+Subsequently, inspired by Zero-shot-CoT Kojima et al.
+(2022), we establish a series of rules, exemplified in Figure 4 and disassemble them to some subtopics to ensure the discriminating accuracy where the LLM-based discriminator can analyze the generation step by step.
+By adopting this method, the discrimination rules can be modified partially to address certain issues.
+After the discrimination process, as shown in Figure 1, each instruction instance is classified as either a good or bad case and the classification information is subsequently random selected in the following generation as examples.
+For the reusage of these classified instruction instance, different from self-instruct Wang et al.
+(2023a) which solely utilize the initial seed task as good example, we exploit both the good generation and bad generation as few-shot example so that the generator can learn from the mistake in different bad example.
+Therefore, this framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality training dataset.
+| Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
+| GPT-4 | 47.0 | 48.2 | 50.0 | 50.6 | 47.6 | 43.3 | 47.8 |
+| StarCoder | 8.7 | 15.7 | 13.3 | 20.1 | 15.6 | 6.7 | 13.4 |
+| OctoCoder | 30.4 | 28.4 | 30.6 | 30.2 | 26.1 | 16.5 | 27.0 |
+| WizardCoder | 31.8 | 29.5 | 30.7 | 30.4 | 18.7 | 13.0 | 25.7 |
+| WaveCoder-SC-15B |
+39.3 | 35.1 | 34.8 | 36.2 | 30.2 | 22.5 | 33.0 |
+| CodeLLaMa-instruct-7B | 28.0 | 23.2 | 23.2 | 18.3 | 0.1 | 0.1 | 15.5 |
+| CodeLLaMa-CodeAlpaca-7B | 37.8 | 39.0 | 42.0 | 37.8 | 37.2 | 29.2 | 37.1 |
+| WaveCoder-CL-7B |
+41.4 | 41.4 | 42.0 | 47.1 | 42.7 | 34.7 | 41.5 |
+| CodeLLaMa-instruct-13B | 29.2 | 19.5 | 32.3 | 24.4 | 12.8 | 0.1 | 19.7 |
+| CodeLLaMa-CodeAlpaca-13B | 42.7 | 43.9 | 50.0 | 45.7 | 39.6 | 37.2 | 43.2 |
+| WaveCoder-CL-13B |
+48.8 | 48.2 | 50.6 | 51.8 | 45.1 | 40.2 | 47.4 |
+| DeepseekCoder-6.7B | 29.9 | 29.2 | 39.0 | 29.2 | 25.0 | 21.9 | 29.0 |
+| Magicoder-DS | 42.0 | 43.3 | 50.6 | 41.4 | 38.4 | 29.2 | 40.8 |
+| DeepseekCoder-CodeAlpaca-6.7B | 49.4 | 51.8 | 45.1 | 48.8 | 44.5 | 31.7 | 45.2 |
+| WaveCoder-DS-6.7B |
+57.9 | 52.4 | 57.3 | 47.5 | 45.1 | 36.0 | 49.4 |
+| WaveCoder-Pro-6.7B |
+59.1 | 56.7 | 54.2 | 45.1 | 45.7 | 34.1 | 49.2 |
+| Deepseek-instruct-6.7B | 56.1 | 58.5 | 57.3 | 49.4 | 45.1 | 36.6 | 50.5 |
+| Magicoder-S-DS | 56.1 | 55.4 | 58.5 | 51.2 | 45.7 | 35.3 | 50.3 |
+| WaveCoder-Ultra-6.7B |
+58.5 | 57.3 | 61.0 | 53.0 | 50.0 | 37.2 | 52.8 |
 3 Experiments
-3.1 Training
-Unlike the previous work [14, 32, 33] that mainly focus on code generation task, we hope that CodeOcean could gain more general problem-solving capabilities on a wider range of coding tasks.
-Therefore, our training relies on a 20k instruction dataset generated by our LLM-based Generator-Discriminator framework and an additional logical dataset.
-The 20k dataset covers 4 common code tasks, including code generation, code translation, code summarization, and code repair.
+3.1 Setup
+Unlike the previous work Luo et al.
+(2024); Shen et al.
+(2023); Gunasekar et al.
+(2023) that mainly focus on code generation task, we generate about 20K dataset covers 4 common code-related tasks to enhance the geralization abilties of Code LLM.
 To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU.
-For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5.
-For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5.
-The sequence length of our each instruction sample is less than 1024.
-Followed Alpaca [12], we set the fine-tuning prompt as follows:
-3.2 Evaluation
-Model | Params | Base Model | InsT Data | HumanEval | MBPP (500) |
-Proprietary Models | |||||
-GPT-4 [1] | - | - | - | 86.6 / 67.0 | - |
-ChatGPT [1] | - | - | - | 73.2 / 48.1 | 52.2 |
-Open-Source Models | |||||
-StarCoder | 15B | - | ✘ | 33.6 | 43.3 |
-OctoCoder [21] | 15B | StarCoder | 13K | 46.2 | 43.5 |
-WizardCoder [14] | 15B | StarCoder | 78K | 57.3 | 51.8 |
-WaveCoder-SC-15B ![]() |
-15B | StarCoder | 20K | 50.5 (+16.9) | 51.0 (+7.4) |
-CodeLLaMa [7] | 7B | - | ✘ | 33.5 | 41.4 |
-CodeLLaMa-instruct [7] | 7B | CodeLLaMa | 14K | 34.8 | 44.4 |
-WaveCoder-CL-7B ![]() |
-7B | CodeLLaMa | 20K | 48.1 (+14.6) | 47.2 (+5.8) |
-CodeLLaMa [7] | 13B | - | ✘ | 36.0 | 47.0 |
-CodeLLaMa-instruct [7] | 13B | CodeLLaMa | 14K | 42.5 | 49.4 |
-WaveCoder-CL-13B ![]() |
-13B | CodeLLaMa | 20K | 55.4 (+19.4) | 49.6 (+2.6) |
-DeepseekCoder [22] | 6.7B | - | ✘ | 49.4 | 60.6 |
-Magicoder-DS[34] | 6.7B | DeepseekCoder | 75K | 66.5 | 60.4 |
-WaveCoder-DS-6.7B ![]() |
-6.7B | DeepseekCoder | 20K | 64.0 (+14.6) | 62.8 (+2.2) |
-3.2.1 Benchmarks and Baselines
+For StarCoder-15B, CodeLLaMa-7B and CodeLLaMa-13B, we set the global batch size to 256 using Tensor Parallel and set the initial learning rate at 2e-5.
+For DeepseekCoder-6.7B, we set the global batch size to 512 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5.
+Benchmarks and Baselines.
+To ensure a thorough assessment of the model’s generalization ability, we score our model on three code benchmarks across different code related tasks: HumanEval Chen et al.
+(2021), MBPP Austin et al.
+(2021) and HumanEvalPack Muennighoff et al.
+(2024), as illustrated in Appendix D.
+Proprietary Models.
+We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4.
+If not reported, we use the results from Octopack Muennighoff et al.
+(2024) or evaluate by ourselves.
+Open Source Models. To ensure an equitable comparison, we opted to select models that have been trained with the similar amount of instruction instances for our comparative analysis.
+SoTA Open Source Models.
+We compared WaveCoder-6.7B with the SoTA open source Code LLM, includes Magicoder-S-DS Wei et al.
+(2023) and DeepseekCoder-instruct-6.7B Wei et al.
+(2023) on a wide range of code-related tasks.
+All the result of SoTA open source models is presented from EvalPlus.
+Liu et al.
+(2023) If not reported, we evaluate it by ourselves.
+| Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
+| GPT-4 | 64.6 | 57.3 | 51.2 | 58.5 | 38.4 | 42.7 | 52.1 |
+| StarCoder | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
+| WizardCoder | 32.5 | 33.0 | 27.4 | 26.7 | 28.2 | 16.9 | 27.5 |
+| OctoCoder | 35.1 | 24.5 | 27.3 | 21.1 | 24.1 | 14.8 | 24.5 |
+| WaveCoder-SC-15B |
+37.1 | 33.3 | 40.5 | 23.3 | 31.8 | 19.3 | 30.8 |
+| CodeLLaMa-instruct-7B | 33.5 | 36.0 | 31.7 | 21.3 | 25.0 | 16.4 | 27.3 |
+| CodeLLaMa-CodeAlpaca-7B | 34.7 | 24.4 | 37.8 | 23.2 | 28.6 | 19.5 | 28.0 |
+| WaveCoder-CL-7B |
+41.4 | 31.7 | 39.0 | 25.0 | 34.1 | 23.2 | 32.4 |
+| CodeLLaMa-instruct-13B | 40.2 | 26.8 | 37.2 | 22.5 | 28.0 | 14.6 | 28.2 |
+| CodeLLaMa-CodeAlpaca-13B | 32.3 | 28.0 | 34.1 | 18.9 | 29.9 | 20.7 | 27.3 |
+| WaveCoder-CL-13B |
+45.7 | 42.0 | 48.2 | 32.3 | 38.4 | 20.7 | 37.9 |
+| DeepseekCoder-6.7B | 43.9 | 40.2 | 37.8 | 29.2 | 34.1 | 22.5 | 34.6 |
+| Deepseek-CodeAlpaca-6.7B | 40.8 | 37.2 | 42.1 | 29.9 | 31.7 | 22.5 | 34.0 |
+| Magicoder-DS | 55.5 | 36.6 | 49.4 | 36.0 | 39.6 | 27.4 | 40.7 |
+| WaveCoder-DS-6.7B |
+48.2 | 47.5 | 49.4 | 32.3 | 48.2 | 22.0 | 41.3 |
+| WaveCoder-Pro-6.7B |
+53.0 | 43.3 | 54.9 | 34.1 | 42.7 | 20.0 | 41.3 |
+| Magicoder-S-DS | 60.3 | 46.3 | 54.3 | 38.4 | 48.1 | 29.2 | 46.1 |
+| Deepseek-instruct-6.7B | 62.2 | 54.3 | 61.0 | 39.6 | 55.5 | 33.5 | 51.0 |
+| WaveCoder-Ultra-6.7B |
+56.7 | 50.0 | 54.3 | 34.8 | 51.2 | 36.6 | 47.3 |
+3.2 Result
+Evaluation on Code Generation Task.
+HumanEval and MBPP are two representative benchmarks for code generation task, as illustrated in Appendix D.
+Table 3 shows the pass@1 score of different LLMs on both benchmarks.
+From the results, We have the following observations:
+1) WaveCoder-Pro-6.7B outperforms other open source models with only 6.7B parameters and 20K instruction data.
+Trained with GPT-4 enhanced CodeSeaXDataset dataset, WaveCoder-Pro-6.7B achieve 72.0% pass@1 and on HumanEval and 63.6% on MBPP, surpassing all open source models but still behind proprietary models and the SoTA open source models.
+2) Refined and diverse instruction data can significantly improve the efficiency of instruction tuning.
+As delineated in Table 3, WaveCoder demonstrates commendable performance, utilizing a dataset comprising merely about 20K Instruction Tuning Data (InsT Data), which positions it on an equal footing with its contemporaries.
+Despite a discernible shortfall in the code generation benchmarks relative to WizardCoder (50.5 vs 57.3) and Magicoder (64.0 vs 66.5), it is imperative to consider the substantial disparity in the volume of training data.
+Moreover, it is observed that WaveCoder-pro-6.7B significantly outperforms Magicoder-DS-6.7B (72.0 vs 66.5), demonstrating the effectiveness of data quality and diversity in instruction tuning.
+Evaluation on Other Code-related Task.
+We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack Muennighoff et al.
+(2024) in Table 4 and Table 5, highlighting the the following salient observations:
+1) WaveCoder models outperform all open source models on other code-related task.
+Building upon Starcoder, our proposed WaveCoder-SC has exhibited exceptional performance, transcending the capabilities of both WizardCoder and OctoCoder as evidenced by the HumanEvalFix (33.0 vs 25.7 vs 27.0) and HumanEvalExplain (30.8 vs 27.5 vs 24.5) benchmarks, which is also shown in other base models.
+Notably, WaveCoder-DS-6.7B achieves 49.4% average pass@1 score on HumanEvalFix and 41.3% on HumanEvalExplain, surpassing all open source models and demonstrating strong generalization capabilities in multi-task scenarios.
+2) The enhancement in data refinement and diversification can markedly bolster the efficacy of instruction tuning in multi-task scenarios.
+Such data refinement, coupled with the categorization of instructions into four code-related tasks, has propelled our models to reach an unforeseen generalization capabilities in various code-related tasks.
+Remarkably, our WaveCoder-DS-6.7B model outperforms GPT-4 (49.4 vs 47.8) on HumanEvalFix, thereby underscoring the potential of smaller models to achieve near-parity with parameter-heavy models when optimized efficiently.
+WaveCoder-Ultra-6.7B.
+Inspired by Magicoder-S-DS-6.7B Wei et al.
+(2023), we combine CodeSeaXDataset with WaveCoder-evol-codealpaca to a 130K dataset.
+Fine-tuned with this combination of two datasets, we obtain WaveCoder-Ultra-6.7B.
+As illustrated in Table 3, 4, 5, WaveCoder-Ultra-6.7B has the state-of-the-art generalization abilities on a wide range of code-related tasks, which highlights the significance of our CodeSeaXDataset dataset again and demonstrates the potential of larger datasets.
+| Model | CG | CS | CT | CR | HumanEval | HumanEval Fix (Avg.) | HumanEval Explain (Avg.) |
+| DeepseekCoder-Base-6.7B | ✘ | ✘ | ✘ | ✘ | 49.4 | 29.0 | 34.6 |
+| WaveCoder-DS-6.7B |
+✔ | ✔ | ✔ | ✔ | 64.0 (+14.6) | 49.4 (+20.4) | 41.3 (+7.3) |
+| -Without Repair | ✔ | ✔ | ✔ | ✘ | 60.9 (-3.1) | 15.7 (-33.7) | 41.2 (-0.1) |
+| -Without Generation | ✘ | ✔ | ✔ | ✔ | 53.6 (-10.4) | 47.4 (-2.0) | 40.5 (-0.8) |
+| -Without Translation | ✔ | ✔ | ✘ | ✔ | 60.9 (-3.1) | 49.3 (-0.1) | 41.6 (+0.3) |
+| -Without Summarization | ✔ | ✘ | ✔ | ✔ | 61.5 (-2.5) | 45.6 (-3.8) | 28.4 (-12.9) |
+4 Ablation and Analysis
+4.1 Ablation of Code-related Tasks
+To explore the relationship between different tasks, we conduct an ablation study about the task type of instruction data.
+Using DeepseekCoder-Base-6.7B as our base model and initial 20K CodeSeaXDataset data as our base dataset, we have the following observations from Table 6:
+1) Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff.
+As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks.
+For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
+2) Different tasks can promote each other so that the model can show a generalization ability.
+From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks.
+For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks.
+Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
+4.2 Discussion about Data Leakage
+In this section, we explore the potential leakage through three instruction datasets about code (i.e.
+Code Alpaca, CodeSeaXDataset, Magicoder-evol-codealpaca).
+To ensure an accurate analysis, we employ SoTA embedding model GTE-Large Li et al.
+(2023b) to encode the canonical code in test benchmarks and all code in training set.
+Subsequently, we find the nearest neighbour in train set for each questions in test benchmark.
+As illustrated in Figure 3, CodeSeaXDataset has the lower average cosine similarity than other datasets.
+Figure 6 in Appendix presents two examples about the data leakage in these training set.
+Moreover, we analyze all benchmarks and notice a serious data leakage issue between HumanEval and Magicoder-evol-codealpaca dataset.
+Therefore, we decontaminate Magicoder-evol-codealpaca for each evaluation problem in HumanEval and obtain WaveCoder-evol-codealpaca.
+As illustrated in Figure 3, WaveCoder-evol-codealpaca has lower similarity than Magicoder-evol-codealpaca.
+5 Related Work
+Instruction Tuning.
+Recent studies, such as FLAN Wei et al.
+(2022), ExT5 Aribandi et al.
+(2022), and FLANT5 Chung et al.
+(2022), have underscored the efficacy of integrating diverse tasks within training process to bolster the adaptability of pre-trained models for downstream tasks.
+Specifically, Flan-PaLM 540B’s Chung et al.
+(2022) instruction-tuning over 1.8K tasks has demonstrated that a widespread and versatile enhanced instruction dataset markedly enhances language model performance.
+InstructGPT Ouyang et al.
+(2022), with its incorporation of premium instruction data crafted by human annotators, has shown significant promise in aligning model outputs with user intents, prompting further investigation into instruction-tuning mechanisms.
+Additionally, Stanford Alpaca Taori et al.
+(2023) has innovatively employed GPT-generated instruction data via self-instruct Wang et al.
+(2023a) for instruction tuning process.
+WizardLM Xu et al.
+(2024) has built upon these advancements by applying the evol-instruct methodology, collectively illuminating the transformative impact of instruction tuning on the overall capabilities of LLM.
+Code Large Language Models.
+Recent advancements in code generation have been propelled by Code LLMs such as CodeGen Nijkamp et al.
+(2022), CodeT5 Wang et al.
+(2021), StarCoder Li et al.
+(2023a), CodeLLaMa Roziere et al.
+(2023) and Deepseek-Coder Guo et al.
+(2024), which benefit from extensive pre-training on expansive code corpora.
+Efforts to further enhance efficiency and problem-solving capabilities have led to the development of instruction-tuned models like InstructCodeT5+ Wang et al.
+(2023b), WizardCoder Luo et al.
+(2024), Pangu-coder2 Shen et al.
+(2023), However, all the instruction data they used is from Code Alpaca, which is not refined enough in the context of multi-task environment, which drives us to propose new methods for instruction data generation.
+Concurrently, with the release of our contemporaneous work Magicoder Wei et al.
+(2023), we offer a concise analysis in Section 3.
+6 Conclusion
+This paper presents WaveCoder, a Code LLM fine-tuned with widespread and versatile enhanced instruction data.
+By enabling language models to effectively tackle complex code-related tasks, our approach demonstrates the potential of integrating multiple code-related tasks into instruction tuning for Code LLM and generating high-quality and diverse instruction data for specific task requirements in multi-task scenarios.
+WaveCoder achieves state-of-the-art generalization performance on different code-related tasks surpassing existing open source Code LLMs.
+Furthermore, our analysis of the relationship of different tasks provides valuable insights for future research, paving the way for more extensive code-related tasks and larger dataset.
+Limitations
+We present WaveCoder and propose a data generation method which can stably generate high-quality and diversity instruction data from open source dataset in multi-task scenario.
+One limitation of our work is that the training dataset we used only includes 19,915 instructions, which produces limited enhancements to the model.
+As illustrated Section 3, we expand the training dataset to a larger amount and the resulted model still have significant improvement.
+Therefore, future work should focus on more code-related task types and larger dataset.
+Ethics Statement
+We constructed our CodeSeaXDataset dataset from open source code.
+For each code snippet we used, we are committed to adhering to the terms of its license, which includes proper attribution ensuring that any modifications or derivative works are also shared under the compatible terms.
+Moreover, we notice a serious data leakage issue in the Magicoder-evol-codealpaca dataset.
+To ensure a fair comparison, we remove three nearest neighbours of each question in test benchmark from train set.
+However, if all similar samples are accidentally removed, the integrity of the data will be damaged, which is harmful to model training.
+Therefore, this phenomenon should be attributed to the fact that the problems in the current test benchmarks are some of the most basic algorithm logic.
+To this end, we call for more comprehensive and complex test benchmarks for Code LLMs which will not easily cause data leakage problem.
+Appendix A Prompt
+Followed Alpaca Taori et al. (2023), we set the fine-tuning prompt as follows:
+Appendix B Examples of the LLM-based Generator-Discriminator framework
+| LLM-based Generator-Discriminator framework |
+| Input: |
+| The Python code calculates the area of a circle given its radius. |
+| Output: |
+| task_name: Calculate Circle Area |
+| instruction: Write a Python function that calculates the area of a circle given its radius. |
+| information: The formula to calculate the area of a circle is A = pi * , where A is the area and r is the radius. |
+| solution: |
+| import math def area_of_circle(radius): return math.pi * radius ** 2 |
+| Analysis: - Step 1: Check the Instruction: The programming language should be specified in the instruction.
+<answer: yes, The instruction mentions "Write a Python function," which clearly specifies the programming language as Python.> |
+| - Step 2: Check the Solution: |
+| 1.
+Solution is related to instruction and information.
+Solution is the specific resolution to instruction and information.
+<answer: yes, The solution provided calculates the area of a circle given its radius, which is directly related to the instruction and the information provided.> |
+| 2.
+In the solution, it should only contain the code and comments within the code.
+There should be no explanations provided outside the code.
+<answer: yes, The solution only contains the code without any explanations outside the code.
+Comments within the code are not present, but the code is straightforward and easy to understand.> |
+| 3.
+The code that contains algorithmic logic in solution should not often be buried inside complex or poorly documented functions, making them difficult to follow or learn from.
+<answer: yes, The code that contains algorithmic logic in the solution should not often be buried inside complex or poorly documented functions, making them difficult to follow or learn from.> |
+| 4.
+The code should import the required necessary libraries.
+<answer: yes, The code imports the necessary math library, which is required for calculating the area of a circle using the mathematical constant pi.> |
+| - Overall answer: yes |
+| - Reasons: All the requirements are met as per the given rules.
+The programming language is specified, and the solution is related to the instruction and information.
+The code is simple, easy to follow, and imports the necessary library.
+The solution provided is a good example of a well-structured and concise code that meets the task’s specifications.
+|
+Appendix C Comparison with CodeAlpaca
+CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct Taori et al.
+(2023).
+To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeSeaXDataset), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts.
+To prevent overfitting, we use Low-Rank Adaption (LoRA) Hu et al.
+(2022) for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset.
+1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 5.
+Taking Starcoder as the base model, CodeSeaXDataset surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 5 (a), which emphasizes the effectiveness of our method on refining instruction data.
+As shown in Figure 5 (b), The results of different base models on CodeSeaXDataset surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeSeaXDataset dataset in enhancing the instruction-following ability of the base model.
+2) According to Table 4 and Table 5, All WaveCoder models significantly outperform the model fine-tuned with CodeAlpaca.
+Remarkably, The pass@1 score of WaveCoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain.
+This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
+Appendix D Evaluation Benchmark
 HumanEval 444https://huggingface.co/datasets/openai_humaneval, which consists of 164 manually-written Python programming problems and an average of 9.6 test cases allocated to each problem is now the most extensively adopted benchmark for Code LLMs.
@@
 In this paper, we select the HumanEvalFix to evaluate the code to code ability especially on code repair task and HumanEvalExplain benchmarks to evaluate the code to text ability especially on code summarization task.
-Proprietary Models.
-We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4.
-If not reported, we use the results from Octopack [21] or evaluate by ourselves.
-Open Source Models.
-Base models comprise StarCoder, CodeLLaMa and Deepseek-Coder with prompting.
-To ensure an equitable comparison, we opted to select those instructed models that have been trained using fewer than 100,000 instruction instances for our comparative analysis.
-3.2.2 metrics
-Pass@k.
-Followed most previous works, we also adopt the pass@k metric proposed by OpenAI [4] to evaluate the performance of our model.
-By generating samples for each promblem and checking the correctness for each sample with the test cases, the pass@k score can be calculated by the unbiased estimator:
-where is the total number of sample generated for a problem, is the number of correct samples passing all the test cases.
-3.2.3 Evaluation on Code Generation Task
-HumanEval and MBPP are two representative benchmarks for code generation task where the model should generate the complete code based on the function signature and the docstring of the problem.
-Table 3 shows the pass@1 score of different LLMs on both benchmarks.
-From the results, We have the following observations:
-WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data) and CodeAlpaca with 20K data.
-Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
-Refined and diverse instruction data can significantly improve the efficiency of instruction tuning.
-According to the comparison in Table 3, WaveCoder solely relies on a dataset encompassing 20,000 entries to accomplish performance levels on par with those of Magicoder [34] and Wizardcoder [14].
-Although WaveCoder falls behind WizardCoder [14] and Magicoder [34] on code generation task, there is a considerable difference in data size among them.
-Nevertheless, the performance lag of WaveCoder is relatively minor.
-Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
-GPT-4 | 47.0 | 48.2 | 50.0 | 50.6 | 47.6 | 43.3 | 47.8 |
-StarCoder | 8.7 | 15.7 | 13.3 | 20.1 | 15.6 | 6.7 | 13.4 |
-OctoCoder | 30.4 | 28.4 | 30.6 | 30.2 | 26.1 | 16.5 | 27.0 |
-WizardCoder | 31.8 | 29.5 | 30.7 | 30.4 | 18.7 | 13.0 | 25.7 |
-WaveCoder-SC-15B ![]() |
-39.3 | 35.1 | 34.8 | 36.2 | 30.2 | 22.5 | 33.0 |
-CodeLLaMa-instruct-7B | 28.0 | 23.2 | 23.2 | 18.3 | 0.1 | 0.1 | 15.5 |
-CodeLLaMa-CodeAlpaca-7B | 37.8 | 39.0 | 42.0 | 37.8 | 37.2 | 29.2 | 37.1 |
-WaveCoder-CL-7B ![]() |
-41.4 | 41.4 | 42.0 | 47.1 | 42.7 | 34.7 | 41.5 |
-CodeLLaMa-instruct-13B | 29.2 | 19.5 | 32.3 | 24.4 | 12.8 | 0.1 | 19.7 |
-CodeLLaMa-CodeAlpaca-13B | 42.7 | 43.9 | 50.0 | 45.7 | 39.6 | 37.2 | 43.2 |
-WaveCoder-CL-13B ![]() |
-48.8 | 48.2 | 50.6 | 51.8 | 45.1 | 40.2 | 47.4 |
-DeepseekCoder-6.7B | 29.9 | 29.2 | 39.0 | 29.2 | 25.0 | 21.9 | 29.0 |
-Magicoder-DS | 42.0 | 43.3 | 50.6 | 41.4 | 38.4 | 29.2 | 40.8 |
-DeepseekCoder-CodeAlpaca-6.7B | 49.4 | 51.8 | 45.1 | 48.8 | 44.5 | 31.7 | 45.2 |
-WaveCoder-DS-6.7B ![]() |
-57.9 | 52.4 | 57.3 | 47.5 | 45.1 | 36.0 | 49.4 |
-3.2.4 Evaluation on Other Code-related Task
-We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4, 5 and use the result reported in Octopack [21] when avaliable.
-Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
-WaveCoder outperforms all open source models on other code-related task with similar degrees of fine-tuning.
-Using Starcoder as the foundation model, our WaveCoder-SC outperforms both WizardCoder and OctoCoder across all programming languages on HumanEvalFix and HumanEvalExplain benchmarks.
-Even when compared with other foundation models, our model consistently demonstrates superior performance over all previously fine-tuned models.
-Enhancing the refinement and diversity of data can significantly improve the effectiveness of instruction tuning in multi-task scenarios.
-By refining the data and classifying the instructions to 4 different code-related downstream tasks, our models achieve a surprising performance level on both evaluation scenarios.
-Notably, WaveCoder-DS, utilizing only 6.7B parameters, approximates the performance of GPT-4 on HumanEvalFix quite closely.
-Model | Python | JavaScript | Java | Go | C++ | Rust | Avg. |
-GPT-4 | 64.6 | 57.3 | 51.2 | 58.5 | 38.4 | 42.7 | 52.1 |
-StarCoder | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
-WizardCoder | 32.5 | 33.0 | 27.4 | 26.7 | 28.2 | 16.9 | 27.5 |
-OctoCoder | 35.1 | 24.5 | 27.3 | 21.1 | 24.1 | 14.8 | 24.5 |
-WaveCoder-SC-15B ![]() |
-37.1 | 33.3 | 40.5 | 23.3 | 31.8 | 19.3 | 30.8 |
-CodeLLaMa-instruct-7B | 33.5 | 36.0 | 31.7 | 21.3 | 25.0 | 16.4 | 27.3 |
-CodeLLaMa-CodeAlpaca-7B | 34.7 | 24.4 | 37.8 | 23.2 | 28.6 | 19.5 | 28.0 |
-WaveCoder-CL-7B ![]() |
-41.4 | 31.7 | 39.0 | 25.0 | 34.1 | 23.2 | 32.4 |
-CodeLLaMa-instruct-13B | 40.2 | 26.8 | 37.2 | 22.5 | 28.0 | 14.6 | 28.2 |
-CodeLLaMa-CodeAlpaca-13B | 32.3 | 28.0 | 34.1 | 18.9 | 29.9 | 20.7 | 27.3 |
-WaveCoder-CL-13B ![]() |
-45.7 | 42.0 | 48.2 | 32.3 | 38.4 | 20.7 | 37.9 |
-DeepseekCoder-6.7B | 43.9 | 40.2 | 37.8 | 29.2 | 34.1 | 22.5 | 34.6 |
-Deepseek-CodeAlpaca-6.7B | 40.8 | 37.2 | 42.1 | 29.9 | 31.7 | 22.5 | 34.0 |
-Magicoder-DS | 55.5 | 36.6 | 49.4 | 36.0 | 39.6 | 27.4 | 40.7 |
-WaveCoder-DS-6.7B ![]() |
-48.2 | 47.5 | 49.4 | 32.3 | 48.2 | 22.0 | 41.3 |
-3.3 Comparison with CodeAlpaca
-CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct [12].
-To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeOcean), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts.
-To prevent overfitting, we use Low-Rank Adaption (LoRA) [35] for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset.
-Table 4, 5 present the result of CodeOcean on HumanEval and other code-related benchmarks highlighting the following salient observations:
-1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1.
-Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data.
-As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
-2) According to Table 4 and Table 5, All WaveCoder models significantly outperform the model fine-tuned with CodeAlpaca.
-Remarkably, The pass@1 score of WaveCoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain.
-This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
-4 Ablation and Analysis
-To evaluate the the effectiveness of each code-related task appearing in training data and the relationship between different tasks, we conduct an ablation study about the task type of instruction data.
-Using DeepseekCoder-Base-6.7B as our base model, we have the following observations from Table 6:
-Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff.
-As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks.
-For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
-Different tasks can promote each other so that the model can show a generalization ability.
-From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks.
-For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks.
-Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
-Model | CG | CS | CT | CR | HumanEval | HumanEval Fix (Avg.) | HumanEval Explain (Avg.) |
-DeepseekCoder-Base-6.7B | ✘ | ✘ | ✘ | ✘ | 49.4 | 29.0 | 34.6 |
-WaveCoder-DS-6.7B ![]() |
-✔ | ✔ | ✔ | ✔ | 64.0 (+14.6) | 49.4 (+20.4) | 41.3 (+7.3) |
--Without Code Repair | ✔ | ✔ | ✔ | ✘ | 60.9 (-3.1) | 15.7 (-33.7) | 41.2 (-0.1) |
--Without Code Generation | ✘ | ✔ | ✔ | ✔ | 53.6 (-10.4) | 47.4 (-2.0) | 40.5 (-0.8) |
--Without Code Translation | ✔ | ✔ | ✘ | ✔ | 60.9 (-3.1) | 49.3 (-0.1) | 41.6 (+0.3) |
--Without Code Summarization | ✔ | ✘ | ✔ | ✔ | 61.5 (-2.5) | 45.6 (-3.8) | 28.4 (-12.9) |
-5 Related Work
-5.1 Code Large Language Models
-Numerous researchers have proposed their Code Large Language Models (LLMs), including CodeGen [36], CodeT5[37], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain.
-By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios.
-In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [38] and WizardCoder [14] have been introduced.
-When drafting this paper, we noticed that our contemporaneous work Magicoder [34] has been released, so we conduct a brief comparative analysis in Section 3.
-5.2 Instruction Tuning
-In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language.
-For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users.
-This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT.
-Flan-PaLM 540B [10] is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model.
-Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training.
-For Code LLMs, InstructCodeT5+ [38], WizardCoder [14], Pangu-coder2 [32], are recent recent models trained with coding instructions.
-However, all the instruction data they used is from CodeAlpaca, which is not refined enough.
-6 Conclusion
-This paper studies the effect of multi-task instruction data on enhancing the generalization ability of Code Large Language Model.
-We introduce CodeOcean, a dataset with 20K instruction instances on four universal code-related tasks.
-This method and dataset enable WaveCoder, which significantly imporove the generalization ability of foundation model on diverse downstream tasks.
-Our WaveCoder has shown the best generalization ability among other open-source models in code repair and code summarization tasks, and can maintain high efficiency on previous code generation benchmarks.
-In order to further improve the mono-task performance and generalization ability of the model, future work should focus on the interplay among different tasks and larger dataset.